# ED4: Explicit Data-level Debiasing for Deepfake Detection

### v1

@@ Abstract
Learning intrinsic bias from limited data has been considered the main reason for the failure of deepfake detection with generalizability. Apart from the discovered content and specific-forgery bias, we reveal a novel spatial bias, where detectors inertly anticipate observing structural forgery clues appearing at the image center, also can lead to the poor generalization of existing methods. We present ED4, a simple and effective strategy, to address aforementioned biases explicitly at the data level in a unified framework rather than implicit disentanglement via network design. In particular, we develop ClockMix to produce facial structure preserved mixtures with arbitrary samples, which allows the detector to learn from an exponentially extended data distribution with much more diverse identities, backgrounds, local manipulation traces, and the co-occurrence of multiple forgery artifacts. We further propose the Adversarial Spatial Consistency Module (AdvSCM) to prevent extracting features with spatial bias, which adversarially generates spatial-inconsistent images and constrains their extracted feature to be consistent. As a model-agnostic debiasing strategy, ED4 is plug-and-play: it can be integrated with various deepfake detectors to obtain significant benefits. We conduct extensive experiments to demonstrate its effectiveness and superiority over existing deepfake detection approaches.
@@
Introduction
The growing threats posed by deepfake technology in social media have heightened the necessity of detecting malicious deepfake content. Hence, deepfake detection technology attracts increasing attention from the research community. Most deepfake detectors perform promisingly when dealing with in-dataset images. However, their effectiveness faces significant challenges when transferred to unseen data distribution, which restricts their ability in practical usage.
To improve the generalization ability, existing deepfake detectors make attempts from multiple aspects (Nguyen, Yamagishi, and Echizen 2019; Haliassos et al. 2021; Huang et al. 2023; Liang, Shi, and Deng 2022; Yan et al. 2023a). Among them, methods targeted at removing the model bias hold the view that, deepfake detectors could easily learn the biased information to categorize input data, rather than digging the intrinsic forgery evidence. Liang et al.(Liang, Shi, and Deng 2022) demonstrate that detectors erroneously learn the identity and background information, which can be referred to as content bias. Then, they design an encoder-decoder network, attempting to achieve feature disentanglement through implicit network constraints. Yan et al.(Yan et al. 2023a) contend that detectors tend to focus on the forgery artifacts related to one specific manipulation method, thereby overlooking common forgery artifacts. To address such specific-forgery bias, they also adopt a similar encoder-decoder network to disentangle common forgery features. With inspiring observations and analyses, these methods struggled to implicitly separate the generalizable forgery features from bias with an encoder-decoder network, which is non-intuitive and insufficient to guarantee successful disentanglement. Broadly speaking, the fake synthesis methods can be considered as a way of removing the model bias by creating new fake images. However, their performance is hindered by the limitation in mixing identities (Shiohara and Yamasaki 2022; Dong et al. 2023), fake regions (Chen et al. 2022), or blend types (Li et al. 2020a).
@@ Introduction
In this paper, we propose Explicit Data-level Debiasing for Deepfake Detection (ED4) aiming to address the aforementioned three biases in a unified framework. Specifically, we introduce two effective modules to remove the model bias in a hybrid manner: ClockMix, and Adversarial Spatial Consistency Module (AdvSCM). ClockMix aims to address the content and specific-forgery biases via the clockwise mixing of different images. It takes multiple images with arbitrary faces as input, and then performs sector-based mosaicing centered on the face, to obtain a mixture containing different backgrounds, identities, and forgery artifacts. Different from existing mixing strategies, ClockMix is carefully designed for face forgery detection with high flexibility and surprising effectiveness. Then, we propose AdvSCM to tackle spatial bias. Specifically, we introduce an adversarial generator to produce shuffled images with larger spatial inconsistency. Subsequently, we enforce the detector to learn consistent forgery feature representations of images with distinct spatial-inconsistent versions. By introducing AdvSCM, the deepfake detector is forced to ignore spatial deviations and pay more attention to informative local regions, thus gaining stronger discriminant power in identifying subtle forgery patterns rather than inertly anticipating the observation of structural forgery clues at image centers.
By explicitly removing the model bias at the data level, ED4 is a superior alternative to the implicit disentanglement achieved by designing networks (Liang, Shi, and Deng 2022; Yan et al. 2023a), as well as to traditional fake synthesis approaches (Chen et al. 2022; Dong et al. 2023; Li et al. 2020a; Shiohara and Yamasaki 2022). The advantages could be summarized as:
@@
Related Works
Deepfake Detection
@@
Deepfake detection aims to classify the authority of an input image, which may be forged by deepfake manipulations. Some approaches focus primarily on specific facial representations, such as lips movement (Haliassos et al. 2021) and action unit consistency (Bai et al. 2023). Meanwhile, various studies are dedicated to developing optimal neural network structures to improve detection performance, such as MesoNet (Afchar et al. 2018), Xception (Chollet 2017), and CapsuleNet (Nguyen, Yamagishi, and Echizen 2019). In the frequency domain, SPSL (Liu et al. 2021) and SRM (Luo et al. 2021) introduce the phase spectrum and high-frequency noises to enhance the forgery information for training. These methods focus on identifying specific vulnerabilities inherent in Deepfake methods, achieving significant success in detecting fake images that exhibit these specific vulnerabilities.
@@
Owing to the limited data variations, certain specific data characteristics may be exclusively found within the data associated with a particular label. Subsequently, the detector may take a shortcut by solely learning the correlation between certain characteristics and the label, neglecting to achieve an in-depth comprehension of forgery features. We refer to this inert mapping as model bias. Dong et al.(Dong et al. 2023) and Huang et al.(Huang et al. 2023) posit that the target face identity used during face swapping remains in the fake face, leading to implicit identity leakage. Liang et al.(Liang, Shi, and Deng 2022) recognize the issue of content bias in deepfake detection, suggesting that networks might mistakenly treat certain identities or backgrounds as either Fake or Real based on learned biases from the training data without learning the actual forgery information. Subsequently, they design an encoder-decoder network to implicitly disentangle the content bias, attempting to obtain a forgery-only feature for detection. Similarly, specific-forgery bias refers to detectors overfitting method-specific-forgery artifacts, thereby exhibiting a lack of sensitivity to common forgery and demonstrating limited generalizability. UCF (Yan et al. 2023a) utilizes indirect constraints and network design to implicitly disentangle common forgery features from specific features, guiding the learning toward generalized forgery information.
@@
While these methods demonstrate the significance of content and specific-forgery biases, the implicit disentanglement achieved by the encoder-decoder network design appeared overly complex and indirect. Namely, the components for decoding are redundant during inference and are limited to indirectly constrain the encoder for the disentanglement during training.
Data Augmentation and Synthesis for Deepfake Detection
Figure 2: Overall framework of the proposed approach.
@@
Data augmentation is widely employed during the training of neural networks to enhance the generalizability of the model, which enriches data diversity by implementing spatial transformations to the training samples. Various traditional data augmentation methods have been demonstrated to exhibit significant impacts on the generalizability of deepfake detectors. For instance, Liang et al.(Liang, Shi, and Deng 2022) attempts to incorporate the feature-level Mixup, while Wang et al.(Wang and Deng 2021) designed a specific version of random erasing (Zhong et al. 2020) based on attention for deepfake detection. DCL (Sun et al. 2022) proposes to leverage groups of augmented images via dual contrastive learning. Additionally, many strategies introduce diverse data synthesis approaches that are tailored to the deepfake detector. Unlike traditional augmentation, these synthesis methods generate new data from the existing training data and with more defined objectives. Specifically, Face X-ray (Li et al. 2020a) simulates blending artifacts by replacing one pristine face with another pristine face with the nearest landmarks. SLADD (Chen et al. 2022) adopts an adversarial training strategy to select harder forgery artifacts and blend them to pristine faces, obtaining more challenging samples to enhance the sensitivity of the detector. FWA (Li and Lyu 2018) and SBI (Shiohara and Yamasaki 2022) simulated the quality inconsistency by replacing a pristine face with a transformed version of itself, forcing detectors to heighten sensitivity to forged information.
@@
The existing data synthesis methods mixed two paired images with specific regions, which limits the diversity of data mixing. In contrast, our method employs a comprehensive random integration of multiple arbitrary faces with any label, which is beneficial to bias removal.
Proposed Method
Overall Data-level Debiasing Framework
@@
To explicitly remove bias at the data level, we propose a unified framework named ED4 with two essential components. Firstly, we propose ClockMix with arbitrary faces to remove content and specific-forgery biases within the samples. Secondly, we introduce the Adversarial Spatial Consistency Module to address the spatial bias. Our method explicitly extends the training distribution with rich compositions of identity, background, manipulation traces, and spatial arrangements, allowing the detector to learn generalizable forgery representations directly and minimize the undesirable effect of the model bias. As the overall architecture shown in Fig. 2, ED4 is a plug-and-play module that can be directly implemented into the backbone feature extractor.
ClockMix
@@
To disrupt the shortcut mapping associated with the specific-forgery, background, and identity, we propose ClockMix to randomly integrate faces and backgrounds of arbitrary images. ClockMix performs sector-based mosaicing centered on the face, which is analogous to the rotation of clock hands. Specifically, ClockMix first requires conducting face alignment to ensure that the center of each face is consistently located at the central point of each image according to the face landmarks. Then, we introduce a “clock hand” ray $r_{h}$ starting at the central point $(\delta_{x},\delta_{y})$ and define the angle between $r_{h}$ and $r_{base}$ as $\rho$, where $r_{base}$ is a baseline ray starting at the central point toward a random direction. The area in each image swept by $r_{h}$ upon rotating to $\rho$ will be replaced by the corresponding area from other images within the same mini-batch. To achieve this effect, we calculate a swept-area matrix $\mathbf{M}$ to record the angles between each $r_{(i,j)}$ and the vertical upward-oriented ray $r_{v}$, where $r_{(i,j)}$ denotes the ray starting at the central point and passing the pixel located at point $(i,j)$. Hence, each element in $\mathbf{M}$ can be calculated as:
@@
$=\mathbf{I}_{a}\odot(\mathbf{M}_{base}>\rho_{1})+\mathbf{I}_{base}\odot(\mathbf{M}_{base}\leq\rho_{1}),$
@@
Compared with existing fake synthesizing methods with particular image pair, face region, and label selection (Chen et al. 2022; Dong et al. 2023; Li et al. 2020a), or limited identity and forgery participation (Shiohara and Yamasaki 2022), our ClockMix performs more thorough synthesis with arbitrary images. The deepfake detector is allowed to see mixtures of different identities and backgrounds in both real and fake samples, and the co-occurrence of multiple manipulation clues in one fake image, leading to better removal of the content and specific-forgery bias.
@@
Superiority over existing mixing-based methods. We notice that the proposed ClockMix may share similarities with popular data augmentation strategies that perform pixel mixing, including Mosaic (Bochkovskiy, Wang, and Liao 2020), Mixup (Zhang et al. 2017), and Cutmix (Yun et al. 2019). As shown in Fig. 3, existing strategies are not suitable for debiasing in the deepfake detection task. Specifically, the superiority of ClockMix can be demonstrated by answering two questions: i) What kind of mix is more appropriate for deepfake detection? ii) How to assign the labels for mixed images? For the first question, Mosaic introduces multiple intact faces into one single sample, hence it disrupts the high-level statistical distribution of the training sample and thus confuses the detectors. Mixup can lead to the overlap of local textures between faces and thus obfuscate the details of forgery artifacts. Cutmix cannot guarantee even mixtures of both face regions and backgrounds from different images simultaneously, thereby reducing its effectiveness in addressing content bias. By contrast, ClockMix conducts sector dividing based on the face center, gaining better control over selecting useful regions and assigning labels. As for the second question, Mosaic is designed for object detection, therefore it is not intuitive to assign a classification label. Mixup and CutMix introduce a mixing label strategy with linear interpolation of
@@
with $\lambda\in\left[0,1\right]$. However, in experiments, we observed that linear interpolation of the one-hot labels actually weakens the detector’s ability to identify subtle forgery traces. This is consistent with the basic intuition: an image should be classified as fake even with a small region of manipulated faces. By assigning the hard labels to fake images, ClockMix can greatly increase the sensitivity of the detector to forgery artifacts. More interestingly, we find that referring to the mixture of real images as real also plays a critical role in improving the model performance. As ClockMix inevitably introduces unappealing seams with different image pieces, the deepfake detector might be prone to rely on the seaming artifacts for determining authenticity. When we perform mixing on real images with the assignment of real labels, the detector can get rid of the impact of unnatural seams, and meanwhile, distinguish real from fake with improved accuracy. Please refer to the Ablation Study for more experimental results and illustrations.
Adversarial Spatial Consistency Module
In this paper, besides content and specific-forgery biases, we have identified a new form of bias, namely spatial bias. This bias refers to the detector’s inclination to inertly anticipate observing structural forgery clues at the image center. We believe this issue adversely affects both the effectiveness and robustness of deepfake detection. In terms of effectiveness, inertly focusing on the central area diminishes the sensitivity of the detector to forgery artifacts. Moreover, the anticipation of structural artifacts leads to limited performance in detecting local forgeries. In terms of robustness, it is evident that variations in preprocessing or camera movement are prevalent in practical applications of deepfake detection. Spatial bias compromises the network robustness against such spatial deviation, as detectors primarily concentrate on the image center, yet spatial deviations displace faces and artifacts away from the focal region.

### v2

@@ Abstract
Learning intrinsic bias from limited data has been considered the main reason for the failure of deepfake detection with generalizability. Apart from the discovered content and specific-forgery bias, we reveal a novel spatial bias, where detectors inertly anticipate observing structural forgery clues appearing at the image center, also can lead to the poor generalization of existing methods. We present ED4, a simple and effective strategy, to address aforementioned biases explicitly at the data level in a unified framework rather than implicit disentanglement via network design. In particular, we develop ClockMix to produce facial structure preserved mixtures with arbitrary samples, which allows the detector to learn from an exponentially extended data distribution with much more diverse identities, backgrounds, local manipulation traces, and the co-occurrence of multiple forgery artifacts. We further propose the Adversarial Spatial Consistency Module (AdvSCM) to prevent extracting features with spatial bias, which adversarially generates spatial-inconsistent images and constrains their extracted feature to be consistent. As a model-agnostic debiasing strategy, ED4 is plug-and-play: it can be integrated with various deepfake detectors to obtain significant benefits. We conduct extensive experiments to demonstrate its effectiveness and superiority over existing deepfake detection approaches. Code is available at https://github.com/beautyremain/ED4.
I Introduction
The growing threats posed by deepfake technology in social media have heightened the necessity of detecting malicious deepfake content. Hence, deepfake detection technology attracts increasing attention from the research community. Most deepfake detectors perform promisingly when dealing with in-dataset images. However, their effectiveness faces significant challenges when transferred to unseen data distributions, which restricts their ability in practical usage.
@@ I Introduction
To improve the generalization ability, existing deepfake detectors make attempts from multiple aspects capsule ; lips ; huang2023implicit ; Disentangle ; ucf ; tip2 ; tip3 . Among them, methods targeted at removing the model bias hold the view that, deepfake detectors could easily learn biased information to categorize input data, rather than digging the intrinsic forgery evidence. Liang et al.Disentangle demonstrate that detectors erroneously learn the identity and background information, which can be referred to as content bias. Then, they design an encoder-decoder network, attempting to achieve feature disentanglement through implicit network constraints. Yan et al.ucf contend that detectors tend to focus on the forgery artifacts related to one specific manipulation method, thereby overlooking common forgery artifacts. To address such specific-forgery bias, they also adopt a similar encoder-decoder network to disentangle common forgery features. With inspiring observations and analyses, these methods struggled to implicitly separate the generalizable forgery features from bias with an encoder-decoder network, which is non-intuitive and insufficient to guarantee successful disentanglement. Broadly speaking, the fake synthesis methods can be considered as a way of removing the model bias by creating new fake images. However, their performance is hindered by the limitation in mixing identities SBI ; leakage , fake regions adv , or blend types xray .
@@
In this paper, we propose Explicit Data-level Debiasing for Deepfake Detection (ED4), aiming to address the aforementioned three biases in a unified framework. Specifically, we introduce two effective modules to remove the model bias in a hybrid manner: ClockMix, and Adversarial Spatial Consistency Module (AdvSCM). ClockMix aims to address the content and specific-forgery biases via the clockwise mixing of different images. It takes multiple images with arbitrary faces as input, and then performs sector-based mosaicing centered on the face, to obtain a mixture containing different backgrounds, identities, and forgery artifacts. Different from existing mixing strategies, ClockMix is carefully designed for face forgery detection with high flexibility and surprising effectiveness. Then, we propose AdvSCM to tackle spatial bias. Specifically, we introduce an adversarial generator to produce shuffled images with larger spatial inconsistency. Subsequently, we enforce the detector to learn consistent forgery feature representations of images with distinct spatially inconsistent versions. By introducing AdvSCM, the deepfake detector is forced to ignore spatial deviations and pay more attention to informative local regions, thus gaining stronger discriminant power in identifying subtle forgery patterns rather than inertly anticipating the observation of structural forgery clues at image centers.
Figure 2: Comparisons between debiasing via implicit network design and explicit data level. The red regions indicate the target part for algorithms’ modifications, whereas the parts in the gray regions remain unaltered by the algorithms. Note that the term “unaltered backbone” refers exclusively to the network architecture, while its parameters are typically optimized during training.
As shown in Fig. 2, the explicit debiasing method pre-processes the training samples instead of altering the backbone network architecture, allowing plug-and-play effectiveness, more intuitive visual understanding, and easier real-world implementation. Therefore, by explicitly removing the model bias at the data level, ED4 is a superior alternative to the implicit disentanglement achieved by designing networks Disentangle ; ucf , as well as to traditional fake synthesis approaches adv ; leakage ; xray ; SBI . The advantages of ED4 could be summarized as:
@@
II Related Works
II-A Deepfake Detection
@@
Deepfake detection aims to classify the authority of an input image, which may be forged by deepfake manipulations. Some approaches focus primarily on specific facial representations, such as lip movement lips and action unit consistency aunet . Meanwhile, various studies are dedicated to developing optimal neural network structures to improve detection performance, such as MesoNet mesonet , Xception xception , and CapsuleNet capsule . In the frequency domain, SPSL spsl and SRM srm introduce the phase spectrum and high-frequency noises to enhance the forgery information for training. IFFD tip1 learns patch-channel correspondence to achieve a more interpretable deepfake detection. These methods focus on identifying specific vulnerabilities inherent in Deepfake methods, achieving significant success in detecting fake images that exhibit these specific vulnerabilities.
@@
Owing to the limited data variations, certain specific data characteristics may be exclusively found within the data associated with a particular label. Subsequently, the detector may take a shortcut by solely learning the correlation between certain characteristics and the label, neglecting to achieve an in-depth comprehension of forgery features. We refer to this inert mapping as model bias. Dong et al.leakage and Huang et al.huang2023implicit posit that the target face identity used during face swapping remains in the fake face, leading to implicit identity leakage. Liang et al.Disentangle design an encoder-decoder network to implicitly disentangle the content bias, attempting to obtain a forgery-only feature for detection. UCF ucf utilizes indirect constraints and network design to implicitly disentangle common forgery features from specific features, guiding the learning toward generalized forgery information.
@@
While these methods demonstrate the significance of content and specific-forgery biases, the implicit disentanglement achieved by the encoder-decoder network design appeared overly complex and indirect. Namely, the components for decoding are redundant during inference and are limited to indirectly constraining the encoder for the disentanglement during training.
II-B Data Augmentation and Synthesis for Deepfake Detection
Figure 3: Overall framework of the proposed approach. Both real and fake training samples are first processed by ClockMix, generating multiple mixed samples according to the mixing combinations. Then, AdvSCM is applied to the mixed samples to remove spatial bias in an adversarial manner. Concatenation represents combining two groups of images together as the new input. Approximation indicates the non-differentiable gradient flow for network optimization is approximated via the Reinforce algorithm, instead of using direct back-propagation. Finally, the ED4 processed samples are input to the backbone detector for supervised learning.
@@ II-B Data Augmentation and Synthesis for Deepfake Detection
Data augmentation is widely employed during the training of neural networks to enhance the generalizability of the model, which enriches data diversity by implementing spatial transformations to the training samples. Various traditional data augmentation methods have been demonstrated to exhibit significant impacts on the generalizability of deepfake detectors. For instance, Liang et al.Disentangle attempts to incorporate the feature-level Mixup, while Wang et al.masking designed a specific version of random erasing erasing based on attention for deepfake detection. DCL dcl proposes to leverage groups of augmented images via dual contrastive learning. TALL tall incorporates consecutive frames in the same video and re-arranges their layout in a thumbnail manner. By doing this, it can calculate self-attention among local patches from different adjacent frames. However, TALL cannot address content bias since frames in the same video share the same content. Additionally, many strategies introduce diverse data synthesis approaches that are tailored to the deepfake detector. Unlike traditional augmentation, these synthesis methods generate new data from the existing training data and with more defined objectives. Specifically, Face X-ray xray simulates blending artifacts by replacing one pristine face with another pristine face with the nearest landmarks. SLADD adv adopts an adversarial training strategy to select harder forgery artifacts and blend them to pristine faces, obtaining more challenging samples to enhance the sensitivity of the detector. FWA FWA and SBI SBI simulated the quality inconsistency by replacing a pristine face with a transformed version of itself, forcing detectors to heighten sensitivity to forged information.
@@
The existing data synthesis methods mix two paired images with specific regions, which limits the diversity of data mixing. In contrast, our method employs a comprehensive random integration of multiple arbitrary faces with any label, which is beneficial to bias removal.
III Analysis of Model Bias
III-A Model Bias and Spurious Correlation
Model bias is a common issue in deep learning and has been widely studied butterfly ; ood-survey ; tench ; domainnet . It is not inherently linked to the network architecture of the model (e.g., CNN, RNN, Transformer), but rather arises from biases present in the training data distribution. Specifically, the network learns spurious correlations between the input samples and the output targets from the data. For instance, most images labeled as “butterfly" in ImageNet also contain flowers butterfly , and most images labeled as “tench" depict a fisherman holding the fish tench . Consequently, the biases present in these data distributions lead the model to inertly establish spurious correlations between certain labels and features that are irrelevant to the task objective sc1 . Finally, the bias makes the networks suffer from Lack of robustness, Lack of reliable confidence estimates, and Suboptimal generalization simple-bias .
III-B In the Context of Deepfake Detection
Here, we illustrate the influence of model bias in deepfake detection. As shown in the first row of Fig. 1, if the training data only contains images labeled as “fake" with striped backgrounds or prominent flashes, the model will tend to erroneously associate those patterns with the "fake" label. This not only causes the detector to make misclassifications due to learned spurious correlations based on certain features, but also reduces the detector’s sensitivity to core forgery clues, thereby further degrading its performance. In this paper, we address three major biases in deepfake detection, that is, content bias, specific forgery bias, and spatial bias.
Content and Specific Forgery Biases. Content and specific forgery biases are introduced by Disentangle and ucf to deepfake detection. Specifically, Liang et al.recognize the issue of content bias in deepfake detection, suggesting that networks might mistakenly treat certain identities or backgrounds as either Fake or Real based on learned biases from the training data without learning the actual forgery information. Yan et al.define specific forgery bias as the network overly focusing on forgery clues that only exist in specific forgery techniques in training data, thereby undermining its generalization ability to unseen forgeries.
The Novel Spatial Bias. In this paper, besides content and specific-forgery biases, we have identified a new form of bias, namely spatial bias. It refers to the detector’s inclination to inertly anticipate observing structural forgery clues at the image center. For example, as shown in the third row of Fig. 1, the model consistently focuses on the image center and ignores the localized and shifted forgery clues. We believe this issue adversely affects both the effectiveness and robustness of deepfake detection.
  • Impact on Effectiveness: inertly focusing on the central area diminishes the sensitivity of the detector to forgery artifacts, which further encourages the model to completely rely on the simple features, while neglecting similarly predictive complex features simple-bias . Moreover, the anticipation of structural artifacts leads to limited performance in detecting local forgeries.
  • Impact on Robustness: it is evident that variations in preprocessing or camera movement are prevalent in practical applications of deepfake detection. Spatial bias compromises the network robustness against such spatial deviation, as detectors primarily concentrate on the image center, yet spatial deviations displace faces and artifacts away from the focal region. Experiments about robustness could be found in Sec. V-E
IV Proposed Method
IV-A Overall Data-level Debiasing Framework
@@
To explicitly remove bias at the data level, we propose a unified framework named ED4 with two essential components. Firstly, we propose ClockMix with arbitrary faces to remove content and specific-forgery biases within the samples. Secondly, we introduce the Adversarial Spatial Consistency Module to address the spatial bias. Our method explicitly extends the training distribution with rich compositions of identity, background, manipulation traces, and spatial arrangements, allowing the detector to learn generalizable forgery representations directly and mitigate the adverse effects of model bias. As the overall architecture in Fig. 3, ED4 is a plug-and-play module that can be directly implemented into the backbone feature extractor. Namely, ED4 can be applied to the training samples without disrupting the original DFD training pipeline of the backbone network. While during the inference stage, our method requires only the ED4-trained feature extractor (same network architecture as the backbone) to perform detection.
IV-B ClockMix
@@
To disrupt the shortcut mapping associated with the specific-forgery, background, and identity, we propose ClockMix to randomly integrate faces and backgrounds of arbitrary images. ClockMix performs sector-based mosaicing centered on the face, which is analogous to the rotation of clock hands. Specifically, ClockMix initially performs face alignment to guarantee that the center of each face is consistently located at the central point of each image according to the face landmarks. Then, we introduce a “clock hand” ray $r_{h}$ starting at the central point $(\delta_{x},\delta_{y})$ and define the angle between $r_{h}$ and $r_{base}$ as $\rho$, where $r_{base}$ is a baseline ray starting at the central point toward a random direction. The area in each image swept by $r_{h}$ upon rotating to $\rho$ will be replaced by the corresponding area from other images within the same mini-batch. To achieve this effect, we calculate a swept-area matrix $\mathbf{M}$ to record the angles between each $r_{(i,j)}$ and the vertical upward-oriented ray $r_{v}$, where $r_{(i,j)}$ denotes the ray starting at the central point and passing the pixel located at point $(i,j)$. Hence, each element in $\mathbf{M}$ can be calculated as:
@@
$=\mathbf{I}_{a}\odot(\mathbf{M}_{base}>\rho_{1})+\mathbf{I}_{b}\odot(\mathbf{M}_{base}\leq\rho_{1}),$
@@
Compared with existing fake synthesizing methods with particular image pair, face region, and label selection adv ; leakage ; xray , or limited identity and forgery participation SBI , our ClockMix performs more thorough synthesis with arbitrary images. The deepfake detector is allowed to see mixtures of different identities and backgrounds in both real and fake samples, and the co-occurrence of multiple manipulation clues in one fake image, leading to better removal of the content and specific-forgery bias.
@@
Superiority over existing mixing-based methods. We observe that the proposed ClockMix may share similarities with popular pixel-wise mixing data augmentation strategies. , including Mosaic yolov4 , Mixup mixup , and Cutmix cutmix . As shown in Fig. 4, existing strategies are not suitable for debiasing in the deepfake detection task. Specifically, the superiority of ClockMix can be demonstrated by answering two questions: i) What kind of mix is more appropriate for deepfake detection? ii) How to assign the labels for mixed images? In the first case, Mosaic introduces multiple intact faces into one single sample, thereby disrupting the high-level statistical distribution of the training data and confusing the detectors. Mixup can lead to the overlap of local textures between faces and thus obfuscate the details of forgery artifacts. Cutmix cannot guarantee even mixtures of both face regions and backgrounds from different images simultaneously, thereby reducing its effectiveness in addressing content bias. By contrast, ClockMix conducts sector dividing based on the face center, gaining better control over selecting useful regions and assigning labels. As for the second question, Mosaic is designed for object detection, therefore it is not intuitive to assign a classification label. Mixup and CutMix introduce a mixing label strategy with linear interpolation of
@@
with $\lambda\in\left[0,1\right]$. However, in experiments, we observed that linear interpolation of the one-hot labels actually weakens the detector’s ability to identify subtle forgery traces. This is consistent with the basic intuition: an image should be classified as fake even with a small region of manipulated faces. By assigning the hard labels to mixed images, ClockMix can greatly increase the sensitivity of the detector to forgery artifacts.
Figure 5: T-SNE t-SNE visualization for feature distributions. O denotes the Original samples and A denotes the ED4-Augmented samples. It could be observed that Ours is Real/Fake separated and A/O are indispensable, while Baseline’s A Real/Fake are distributed next to O-Real. This indicates that the seaming artifacts introduced by our method are ignored by the network, where the distinguished fake and real feature distributions each comprise ClockMix-synthesized and real-world samples of their respective class.
More interestingly, we find that referring to the mixture of real images as real also plays a critical role in improving the model performance. As ClockMix inevitably introduces unappealing seams with different image pieces, the deepfake detector might be prone to relying on the seam artifacts for determining authenticity. When we perform mixing on real images with the assignment of real labels, along with introducing a portion of original samples into training, the detector can get rid of the impact of unnatural seams, and meanwhile, distinguish real from fake with improved accuracy. In Fig. 5, we conduct experiments on both original samples (i.e., O-Real and O-Fake) and augmented samples (i.e., A-Real and A-Fake with seaming artifacts), and visualize the feature distributions of baseline and our method. The results show that our detector could distinguish both O-Real/Fake and A-Real/Fake without splitting each other, demonstrating that the seaming artifacts are ignored by our feature extractor.
In summary, the advantages of ClockMix among existing mixing-based augmentations are twofold:
  • Regionally Controlled Sector-Based Mixing. ClockMix partitions each image into angular sectors centered on the facial midpoint, enabling targeted extraction and recombination of both facial and background regions. This approach preserves critical forgery artifacts while avoiding the high-level distribution distortion introduced by Mosaic, the texture blurring of Mixup, and the unbalanced region sampling of CutMix.
  • Hard-Label Assignment with Real-Data Regularization. By assigning binary “real” or “fake” labels to mixed samples – rather than employing linear interpolation of one-hot vectors – ClockMix maintains maximal sensitivity to subtle manipulation traces. Furthermore, labeling real–real mixtures as genuine regularizes the detector against seam-based artifacts, thereby enhancing robustness and overall classification accuracy. Furthermore, this labeling strategy allows conducting ClockMix to mix multiple images instead of two images, which enhances the diversity of mixing results.
Please refer to the Ablation Study for more experimental results and illustrations.
IV-C Adversarial Spatial Consistency Module